# Hydrodynamics without Boost-Invariance from Kinetic Theory – From Perfect Fluids to Active Flocks

### v2

@@ Appendix D On Derivative Expansions and Stochastic Fluctuations
where $\square_{x}$ is the d’Alembertian operator with respect to the $x$ coordinate (to be compared with (38) and (41) with $-iG^{ij}_{F}$ identified with $K^{ij}$ and with $\kappa$ identified as cPlanck-$\hbar c$, notwithstanding the different dimensions in which these expressions are defined).
@@
$V(r)=\frac{e_{0}^{2}}{4\pi r}\left[1+\frac{\alpha_{0}}{3\pi}\left(\ln\frac{\hbar^{2}}{m_{e}^{2}c^{2}r^{2}}-2\gamma_{E}-\frac{5}{3}\right)+...\right],$ (116)
@@ Appendix E Physical Interpretation of the Equation of Kinetic State
Modeling the momentum exchange of a particle or boid with its environment simplifies greatly if one assumes that any effective drag force from momentum exchange is collinear to the applied force (i.e. one can ignore ‘magnetic’ or curl type effective interactions). Hence, presuming the applied force to be along any arbitrary fixed direction ${\mathbf{F}}\equiv F\,\hat{\bf i}$, so that ${\mathbf{F}}_{\rm drag}$ can also be expressed as ${\mathbf{F}}_{\rm drag}=-\zeta(v)\,\hat{\bf i}$, where $v$ is defined via ${\mathbf{v}}=v\,\hat{\bf i}$, one obtains the applied force equation and corresponding effective inertial mass for a particle of mass $m$ to be:
$m\dot{v}=F-\zeta(v),\qquad m_{\rm in}=\frac{1}{1-\frac{\zeta(v)}{F}},$ (119)
@@
$\partial_{t}\rho+({\mathbf{v}}\cdot\nabla)\rho=g\rho,$ (121)

### v3

@@ Appendix D On Derivative Expansions and Stochastic Fluctuations
where $\square_{x}$ is the d’Alembertian operator with respect to the $x$ coordinate (to be compared with (38) and (41) with $iG^{ij}_{F}$ identified with $K^{ij}$ and with $\kappa$ identified as $\hbar c$, notwithstanding the different dimensions in which these expressions are defined).
@@
[…]
@@ Appendix E Physical Interpretation of the Equation of Kinetic State
Modeling the momentum exchange of a particle or boid with its environment simplifies greatly if one assumes that any effective drag force from momentum exchange is collinear to the applied force (i.e. one can ignore ‘magnetic’ or curl type effective interactions). Hence, presuming the applied force to be along any arbitrary fixed direction $^{\mathbf{F}}\equiv F\,\hat{\bf i}$, so that ${\mathbf{F}}_{\rm drag}$ can also be expressed as $^{\mathbf{F}}_{\rm drag}=-\zeta(v)\,\hat{\bf i}$, where $v$ is defined via $^{\mathbf{v}}=v\,\hat{\bf i}$, one obtains the applied force equation and corresponding effective inertial mass for a particle of mass $m$ to be:
$m\dot{v}=F-\zeta(v),\qquad m_{\rm in}=\frac{m}{1-\frac{\zeta(v)}{F}},$ (122)
@@
$\partial_{t}\rho+\nabla\cdot(\rho{\mathbf{v}})=g\rho,$ (124)